OpenStudy (anonymous):

help.... hi can someone pls. explain to me about Inverse trigonometric functions(Differential Calculus). pls. solve this problem then explain pls. y=square root Arcsinx y=xArcsin2x

OpenStudy (anonymous):

\[y=\sqrt{Arcsin x}\]

myininaya (myininaya):

are you looking for the intersection of the two curves? i don't understand what the question is? what i'm solving for?

OpenStudy (anonymous):

the derivative of the inverse trig. functions

myininaya (myininaya):

so you have two problems? this is not one problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you pls. explain?

myininaya (myininaya):

explanation in the both of pdfs

myininaya (myininaya):

you have a specific question on what i did let me know

OpenStudy (anonymous):

in the part getting the derivative i'm confused

myininaya (myininaya):

you do know the chain rule right?

myininaya (myininaya):

quotient rule?

OpenStudy (anonymous):

yes

myininaya (myininaya):

derivative of x with respect to x is 1 derivative of sin(y^2) =(y^2)'cos(y^2)=2yy'cos(y^2)

OpenStudy (anonymous):

chain rule dy/dx=dy/du*du/dx right?

OpenStudy (anonymous):

dang i could have sword i just did these!

myininaya (myininaya):

derivative of 2x=2 derivative of sin(y/x)=(y/x)'cos(y/x)=([y'x-y]/x^2)cos(y/x)

myininaya (myininaya):

satellite give her your formula

OpenStudy (anonymous):

hello myininaya

myininaya (myininaya):

hey

myininaya (myininaya):

i don't think she likes way

OpenStudy (anonymous):

all written out, but maybe a step or two is not clear i don't know.

myininaya (myininaya):

my way*

OpenStudy (anonymous):

\[\frac{d}{dx}\sqrt{f(x)}=\frac{f'(x)}{2\sqrt{f(x)}}\] via the chain rule

OpenStudy (anonymous):

actually i wrote it out, so maybe it is clear, maybe not. you need two facts \[\frac{d}{dx}\arcsin(x)=\frac{1}{\sqrt{1-x^2}}\] and the chain rule to finish

OpenStudy (anonymous):

of course \[\frac{d}{dx}\arcsin(f(x))=\frac{f'(x)}{\sqrt{1-f^2(x)}}\] also by the chain rule. gotta love that chain rule

OpenStudy (anonymous):

thank you myininaya and satellite73